OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (-8, 0) and (1, 5)

OpenStudy (anonymous):

Slope is 5/9 and the y-intercept is 40/9

OpenStudy (anonymous):

@sleepyjess @hartnn

OpenStudy (sleepyjess):

Ok, so we put that into slope intercept form, y = 5/9x + 40/9 What do you think we do now?

OpenStudy (anonymous):

Multiply everything by 9..?

OpenStudy (sleepyjess):

Yes!

OpenStudy (anonymous):

9y=5x+40

OpenStudy (sleepyjess):

Yep, now to get x to the left side

OpenStudy (anonymous):

-5x+9y=40 but don't I need to divide everything by -1?

OpenStudy (sleepyjess):

multiply :)

OpenStudy (anonymous):

Making it 5x-9y=-40

OpenStudy (anonymous):

Bring 9y to right side and +40 to left side..

OpenStudy (sleepyjess):

yep

OpenStudy (anonymous):

Yes yes my bad lol. Blonde moment.

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

THANK YOU AGAIN @sleepyjess :D

OpenStudy (sleepyjess):

We all have them haha